OpenStudy (anonymous):

Can you use the Law of Sines if you are given a = 31, b = 52 and the measure of angle B = 56 degrees? Why or why not?

hero (hero):

From law of sines, you can use the proportion \(\dfrac{\sin A}{a} = \dfrac{\sin B}{b}\) which would allow you to find angle A. Once you find angle A, you would be able to find angle C because \(\angle A + \angle B + \angle C = 180^{\circ}\). From there you would be able to find the length of side c using \(\dfrac{\sin C}{c}\).